OpenStudy (anonymous):

A circle has its center at the center of a square with 3-inch sides. Find the area of the square not covered by the circle Round the answer to the nearest tenth. A) 21.4 in² B) 1.9 in² C) 0.8 in² D) 7.2 in²

OpenStudy (anonymous):

Area of square = 3*3 = 9 inches squared Area of circle = pi*r^2 = pi*(1.5)^2 = 2.25pi inches squared Area not covered by circle = Area of square - Area of circle = 9 - 2.25pi = 1.9 inches squared => B Is this right?

OpenStudy (anonymous):

yup its b

OpenStudy (anonymous):

right